OpenStudy (anonymous):

Find the surface area of a sphere with a radius of 4 inches. Use 3.14 to approximate π. Round answers to the nearest hundredth.

OpenStudy (jdoe0001):

surface area of a sphere = \(\large \boldsymbol{4\pi r^2}\)

OpenStudy (anonymous):

x=4πr^2 x=4*3.14*4^2 x=2524.0576 so it's 2524.06